OpenStudy (anonymous):

what is the x and y intercept of (x^2-3x+2)/(x^2+x-6)

OpenStudy (anonymous):

F(x) = f(x)/g(x) = (x^2 - 3x + 2)/(x^2 + x - 6) F(x) = 0 when f(x) = 0: there are 2 x-intercepts: x = 1 and x = 2. To find y-intercepts, make x = 0 -> y-Intercept F(0) = 2/-6 = -1/3

OpenStudy (anonymous):

thank you @thu1935 really helped me!!